# On-device Anomaly Detection in Conveyor Belt Operations

### v1

@@ Abstract
Mining 4.0 leverages advancements in automation, digitalization, and interconnected technologies from Industry 4.0 to address the unique challenges of the mining sector, enhancing efficiency, safety, and sustainability. Conveyor belts are crucial in mining operations by enabling the continuous and efficient movement of bulk materials over long distances, which directly impacts productivity. While detecting anomalies in specific conveyor belt components, such as idlers, pulleys, and belt surfaces, has been widely studied, identifying the root causes of these failures remains critical due to factors like changing production conditions and operator errors. Continuous monitoring of mining conveyor belt work cycles for anomaly detection is still at an early stage and requires robust solutions. This study proposes two distinctive pattern recognition approaches for real-time anomaly detection in the operational cycles of mining conveyor belts, combining feature extraction, threshold-based cycle detection, and tiny machine-learning classification. Both approaches outperformed a state-of-the-art technique on two datasets for duty cycle classification in terms of F1-scores. The first approach, with 97.3% and 80.2% for normal and abnormal cycles, respectively, reaches the highest performance in the first dataset while the second approach excels on the second dataset, scoring 91.3% and 67.9%. Implemented on two low-power microcontrollers, the methods demonstrated efficient, real-time operation with energy consumption of 13.3 and 20.6 $\mu$J during inference. These results offer valuable insights for detecting mechanical failure sources, supporting targeted preventive maintenance, and optimizing production cycles.
@@ I Introduction
The advancement in Industry 4.0 has transformed industrial operations, redefining manufacturing and production processes by harnessing the power of automation and digitalization. Industry 4.0 creates a framework where intelligent, connected devices enable data sharing and decision-making across processes, strategically integrating cutting-edge technologies such as the Internet of Things, machine learning (ML), big data analytics, and cyber-physical systems [1]. Within this broader context, Mining 4.0 emerges as a specialized domain tailored to the unique challenges of the mining sector. It leverages advanced technologies to streamline exploration, extraction, processing, and transportation processes [2].
@@
Mining companies also invest in these technologies as part of their commitment to safety, sustainability, and resource efficiency. By reducing their environmental footprint, minimizing energy and water waste, and adopting eco-friendly practices, these companies aim to decrease operational risks, ensure compliance with environmental regulations, and align with global sustainability goals [3].
@@
Mining conveyor belts are vital for transporting materials over long distances and rough terrain. These belts allow for continuous and efficient movement of bulk materials, playing a critical role in optimizing productivity and minimizing downtime [4]. The reliable operation of conveyor belts is paramount, as unexpected failures can lead to significant production interruptions, safety risks, and substantial economic losses. Therefore, detecting faults and anomalies in conveyor systems is essential, as it can reduce unexpected downtimes, prevent further damage, and maintain efficient operational workflows [5].
@@
Conveyor belt systems commonly feature various sensors to monitor their operational status and detect early warning signs of mechanical issues. Researchers have developed numerous methodologies using classic ML and deep learning to detect specific faults, such as idler and pulley defects [6], belt damages [7], reduced belt thickness [8], and belt deviation [9]. These approaches leverage data from vibration, thermal, X-ray, acoustic, radio-frequency, and camera sensors to identify these critical anomalies [10]. While considerable attention has been paid to detecting failures in individual components, finding the root causes of failures is of utmost importance. Failures may occur due to changing production conditions, incorrect control settings, operator errors, or equipment malfunctions, emphasizing the need for robust, continuous operation monitoring of conveyor belts [11].
@@
In various industrial applications, such as predictive maintenance, manufacturing, infrastructure, and energy management, time-series data are employed to detect anomalies in operational productive processes using techniques like statistical analysis, data trend evaluation, and process mining [12, 13]. However, many existing methods are computationally intensive or require complex, costly sensor installations, limiting their feasibility in dynamic real-world environments. Due to these constraints, detecting anomalies in the mining sector remains especially challenging.
Regarding continuous monitoring of mining conveyor belts using time-series data, Chiong et al. [14] introduced unsupervised clustering models to detect anomalies in the cycle time of pumps responsible for greasing machine parts, while Cheng et al. [15] analyzed communication traffic data from a coal mine system to identify operational anomalies and potential cyberattacks. Anomaly deviations of a short belt were studied in controlled laboratory conditions with inertial measurement units using a small dataset [16]. Recently, Martinez-Rau et al. [17] developed a two-stage pattern recognition system for real-time anomaly detection in conveyor belts, using a TinyML classifier to identify internal operational modes [18], followed by pattern-matching rules to classify duty cycles. Although effective in low-power microcontrollers (MCUs), this method relies on accurately classifying internal operation modes and has not been tested for robustness under different environmental conditions. This work is an extension of the former study by presenting two alternative pattern recognition approaches for real-time anomaly detection on low-power MCUs. The first approach uses belt speed to mark the start and end of duty cycles, classifying them through pattern-matching techniques. The second approach employs a TinyML classifier trained to identify anomalous cycles while accounting for possible misclassifications in internal operation modes.
@@
The main contributions are the following: (i) It presents and evaluates two low-computational pattern recognition approaches for identifying abnormal conveyor belt usage in real-time on resource-constrained devices. (ii) The approaches are tested for generalization on an additional dataset. (iii) The approaches are optimized using quantization and implemented on two low-power MCUs, measuring power consumption, memory usage, and inference time.
@@ II Conveyor Belt Operation
A hydraulic conveyor belt system operates by harnessing fluid power to drive the movement of the belt through hydraulic components. It is a highly efficient continuous transport mechanism, capable of moving heavy loads over long distances. In this system, hydraulic charge pumps pressurize fluid, which is directed through control valves. These valves regulate the flow and direction of oil to drive the hydraulic motors attached to the drive pulleys or rollers of the conveyor belt. The resulting rotational motion propels the belt, facilitating the seamless transportation of materials along the designated path. The control system provides flexibility in adjusting the speed and direction of the belt, offering precise control over its operation. Moreover, sensors connected to the control system continuously allow the tracking and logging of key hydraulic parameters.
@@ III-A Datasets
This study utilized two datasets, each derived from parameters monitored by the control system of a conveyor belt located in Kiruna, Sweden. An incremental encoder with a pulsed output was employed to measure the number of motor revolutions per minute (rpm). Additionally, pressure sensors recorded both the high- and low-pressure of the hydraulic motor. These sensors were integrated into the control system through a 4-20 mA interface. The collected sensor data underwent preprocessing to provide a sampling rate of one measurement per minute retrieved from the control system’s user interface. For each sensor, the one-minute interval values represent the average of all readings collected at the sensor’s nominal sampling frequency.
The first dataset, namely DS1, comprises data from the entire months of June 2021, October 2021, January 2022, and April 2022, each representing one of the four seasons. This temporal coverage allowed for a broad range of external temperatures, ranging from -26.1 to 26.0 ºC, significantly impacting oil viscosity in the conveyor system and, consequently, the observed pressure levels. The second dataset, namely DS2, was collected from the same conveyor belt during the whole months of June, August, Oktober, and December 2023 with external temperature between -29.0 to 26.0 ºC. Figure 1 compares DS1 and DS2, showing similar high-pressure and speed values distributions. Low-pressure values of DS2 are different than in DS1 covering a bigger value range with noteworthy differences in the quartiles. This difference could be caused by machine wear and maintenance.
DS1 contains approximately 170,000 one-minute state samples, labeled according to operational states: 10,566 samples (6.2%) correspond to the Off state, 123,225 (72.7%) to the Idle state, 11,454 (6.8%) to the Operational state, and 24,299 (14.3%) to the Active state. Additionally, 603 duty cycles were labeled, with 496 cycles (82.3%) categorized as normal and 107 cycles (17.7%) as abnormal. DS2 contains 751 and 194 normal and abnormal duty cycles. DS1 was used to develop and evaluate the proposed approaches, while DS2 was used to assess the generalization and robustness of the approaches in a different dataset.
The start and end timestamps, along with the ground truth labels for the operation modes and duty cycles, were determined by analyzing the waveform patterns using the Imagimob Studio software (Imagimob AB, Stockholm, Sweden) in collaboration with hydraulic system experts.
@@ III-B Algorithms
This study explores a state-of-the-art and two novel approaches to identify anomalies in conveyor belt operations. The approaches classify the duty cycle usages of a conveyor belt applying pattern recognition techniques.
@@
    Approach-1 proposed by Martinez-Rau et al. [17], consists of four consecutive stages (Fig. 2(a)) and serves as a baseline. The algorithm takes as input the three 1-min sensed variables (motor shaft rotational speed, and the high- and low-pressure of the hydraulic motor), which are used to extract a set of 12 features. Four features consist of the three input values and the differential pressure between the high- and low-pressure. The other eight features are the former four features’ third- and fifth-order moving averages, effectively capturing fast and slow variations in the input variables. The complete feature set feeds an ML model to classify the operation modes labeled every 1 min. Since variations in the operation modes typically occur at lower frequencies, a third-order median filter smooths the predicted labels to reduce label fragmentation. Finally, transitions in the filtered state labels are analyzed through pattern matching to detect and classify normal and abnormal duty cycles.
@@
    Approach-2 offers an alternative method for detecting duty cycles (Fig. 2(b)). Unlike Approach-1, which relies on accurately classifying the operation mode labels to detect and classify duty cycles, Approach-2 bypasses the state classification step and focuses directly on the input speed values. As shown by the histogram in Fig. 1, the speed values tend to cluster around 0 and 45 rpm. Based on this distribution, an empirical threshold of 5 rpm is chosen to determine the boundaries of the duty cycles. Once these boundaries are identified, the remaining steps in Approach-2 follow a similar process to those in Approach-1.
@@
    Approach-3 builds on the concept of Approach-2 but introduces an ML model to classify duty cycles instead of relying on the pattern-matching rules used in the previous approaches (Fig. 2(c)). By employing an ML-based classifier, Approach-3 offers the advantage of generating smoother decision boundaries, which can help mitigate the impact of potential misclassifications of the operation modes.
The operation mode and the duty-cycle classifiers are evaluated with different supervised ML algorithms: decision tree (DT), random forest (RF), extra trees (ET) [19], extreme gradient boosting (XGB) [20], Gaussian naive Bayes (NB), and multi-layer perceptron (MLP) neural network [21].
TABLE I: Main properties of the MCU platforms
MCU CPU core Test Board CPU frequency SRAM Flash FPU Current consumption Voltage
@@
III-C Experimental Setup
The experiments were implemented in Python 3.8.10, utilizing the sklearn 1.2.2 package for tree-based and NB algorithms, and TensorFlow 2.13.0 for the MLP. A grid search was performed for hyperparameter tuning to optimize model performance during training. The maximum depth of the DT was set with a restriction of 10, 20, 30, 40, and 50 splits, as well as without restriction. In all cases, the Gini impurity and the Shannon entropy were considered as criteria for the quality split [21]. For the tree-based ensemble algorithms, the number of trees was set to 10, 25, and 50, with maximum tree depths of 4, 6, 8, and 10. The MLP architecture employed a single hidden layer, with the number of neurons varying between 4 and 15. Training was conducted using stochastic gradient descent, with learning rates of 0.1, 0.01, and 0.001.
@@ III-C Experimental Setup
The proposed three approaches for classifying duty cycles were trained and evaluated in the first experiment using DS1. The operation mode classifiers were trained and tested using a leave-one-month-out cross-validation strategy. In each iteration, three months of data were used for training and hyperparameter optimization, while the remaining month served for testing. To address the class imbalance, synthetic oversampling, and random undersampling techniques were applied to the minority and majority operation mode classes before training [24].
In Approach-1 and Approach-2, the predicted operation mode labels were used to detect and classify duty cycles based on heuristic rules. Similar to the operation mode classifier, the duty cycle classifier in Approach-3 was trained and evaluated using leave-one-month-out cross-validation with balanced training data and hyperparameter optimization. The operation mode classifier was trained first, then the duty cycle classifier was trained and evaluated using the predicted operation mode labels across the four months. This approach allows the duty cycle classifier to learn and adapt to potential misclassifications made by the operation mode classifier.
The generalization capabilities of the proposed approaches were evaluated in the second experiment, where the operation mode and duty cycle classifiers were trained on the entire DS1 dataset and tested on DS2. During training, hyperparameter optimization and class balancing techniques were applied, as in the first experiment.
A key factor to consider was the real-time operational requirements for anomaly monitoring in conveyor belts. To assess resource-constrained hardware requirements, the best-performing models in each dataset were implemented in C code. The best-performing ML models from the previous two experiments were converted to C code using the emlearn 0.19.3 package [^1] https://github.com/emlearn/emlearn/tree/0.19.3 for tree-based and NB models, and LiteRT [^2] https://ai.google.dev/edge/litert for the MLP. Similar to the second experiment, the models were trained on DS1 and tested on DS2. The energy consumption, memory usage, and algorithm performance were analyzed using ML models with 32-bit float resolutions and post-training 8-bit integer full model quantization. The compiled C code, using arm-none-eabi-g++, was deployed on two MCU boards: the Raspberry Pi Pico and the Arduino Nano Sense 33 BLE Lite. Table I provides the specifications of these boards.
@@
where $TP_{i}$ counts the number of correctly classified instances of class $i$, $FP_{i}$ counts the number of recognized instances of class $i$ that do not exist in the ground truth (insertions), and $FN_{i}$ counts the number of ground truth instances of class $i$ that are not recognized (deletions). An overall micro-average F1-score was also computed to aggregate the total false positives, false negatives, and true positives across both classes [26]. Except for the deployment on MCUs, each train/test experiment was repeated with different random seeds ten times to ensure robustness. For Approach-3, 100 iterations were performed by combining the ten repetitions of the duty cycle classifiers with the ten repetitions of the operation mode classifiers. The final metrics reported are the averages over all repetitions.

### v2

@@ Abstract
Conveyor belts are crucial in mining operations by enabling the continuous and efficient movement of bulk materials over long distances, which directly impacts productivity. While detecting anomalies in specific conveyor belt components has been widely studied, identifying the root causes of these failures, such as changing production conditions and operator errors, remains critical. Continuous monitoring of mining conveyor belt work cycles is still at an early stage and requires robust solutions. Recently, an anomaly detection method for duty cycle operations of a mining conveyor belt has been proposed. Based on its limited performance and unevaluated long-term proper operation, this study proposes two novel methods for classifying normal and abnormal duty cycles. The proposed approaches are pattern recognition systems that make use of threshold-based duty-cycle detection mechanisms, manually extracted features, pattern-matching, and supervised tiny machine learning models. The explored low-computational models include decision tree, random forest, extra trees, extreme gradient boosting, Gaussian naive Bayes, and multi-layer perceptron. A comprehensive evaluation of the former and proposed approaches is carried out on two datasets. Both proposed methods outperform the former method, with the best-performing approach being dataset-dependent. The heuristic rule-based approach achieves the highest performance in the same dataset used for algorithm training, with 97.3% for normal cycles and 80.2% for abnormal cycles. The ML-based approach performs better on a dataset including the effects of machine aging, scoring 91.3% for normal cycles and 67.9% for abnormal cycles. Implemented on two low-power microcontrollers, the methods demonstrate efficient, real-time operation with energy consumption of 13.3 and 20.6 $\mu$J during inference. These results offer valuable insights for detecting mechanical failure sources, supporting targeted preventive maintenance, and optimizing production cycles.
@@ I Introduction
The advancement in Industry 4.0 has transformed industrial operations, redefining manufacturing and production processes by harnessing the power of automation and digitalization. This evolution takes the name of Mining 4.0 in the specialized mining sector. Mining 4.0 strategically integrates cutting-edge technologies such as the Internet of Things, machine learning (ML), big data analytics, and cyber-physical systems [1] to improve exploration, extraction, processing, and transportation [2]. Mining companies also invest in these tools as part of their commitment to safety, sustainability, and resource efficiency. An aim is to reduce their environmental footprint by minimizing energy and water waste and adopting eco-friendly practices. By doing this, these companies aim to decrease operational risks, ensure compliance with environmental regulations, and align with global sustainability goals [3].
@@
Mining conveyor belts are vital for transporting materials over long distances and rough terrain. These belts allow for continuous and efficient movement of bulk materials, playing a critical role in optimizing productivity and minimizing downtime [4]. The reliable operation of conveyor belts is paramount, as unexpected failures can lead to significant production interruptions, safety risks, and substantial economic losses. Therefore, detecting faults and anomalies in conveyor systems is essential. It helps to reduce unexpected downtime, prevent further damage, and maintain efficient operational workflows [5].
@@
Conveyor belt systems commonly feature various sensors to monitor their operational status and detect early warning signs of mechanical issues. Researchers have developed numerous methodologies using classic ML and deep learning to detect specific faults, such as idler and pulley defects [6], belt damages [7], reduced belt thickness [8], and belt deviation [9]. These approaches leverage data from vibration, thermal, X-ray, acoustic, radio-frequency, and camera sensors to identify these critical anomalies [10]. While considerable attention has been paid to detecting failures in individual components, finding the root causes of failures is of utmost importance. Failures may occur due to changing production conditions, incorrect control settings, operator errors, or equipment malfunctions. This highlights the need for robust, continuous operation monitoring of conveyor belts [11].
@@
In various industrial applications, such as predictive maintenance, manufacturing, infrastructure, and energy management, time-series data are employed to detect anomalies in operational productive processes using techniques like statistical analysis, data trend evaluation, and process mining [12, 13]. However, many existing methods are computationally intensive or require complex and costly sensor installations, which limits their feasibility in dynamic real-world environments. Due to these constraints, detecting anomalies in the mining sector remains especially challenging.
For continuous conveyor belt monitoring, Chiong et al. [14] applied unsupervised clustering to detect anomalies in pump cycle times, while Cheng et al. [15] analyzed coal mine communication data to identify operational anomalies and cyber threats. Anomaly deviations in a short belt were also studied under controlled laboratory conditions using inertial measurement units and a small dataset [16]. Recently, Martinez-Rau et al. [17] developed a method for identifying normal and abnormal duty cycle operations in an industrial mining conveyor belt. However, this method demonstrates challenges in recognizing abnormal cycles and lacks an in-depth evaluation of machine drift over long periods of time, associated with wear or maintenance. Furthermore, while the method demonstrated real-time operation on low-power microcontrollers (MCUs), its energy consumption was only estimated theoretically.
This work is an extension of the former study by comparing the existing approach with two alternative approaches. The proposed methods are lightweight pattern-recognition systems that leverage manually extracted feature sets and a TinyML classifier [18] to identify the internal operation modes. In the first approach, pattern-matching rules analyze the sequence of internal operation modes to classify duty cycles as normal or abnormal. The second approach replaces these heuristic rules with another TinyML classifier trained to detect anomalous cycles while accounting for potential misclassifications in internal operation modes. Six supervised classifiers are evaluated: decision tree (DT), random forest (RF), extra trees (ET) [19], extreme gradient boosting (XGB) [20], Gaussian naive Bayes (NB), and multi-layer perceptron (MLP) neural network [21]. The best-performing models are deployed on two low-power MCUs for real-time anomaly detection.
@@
The main contributions are the following: (i) The study presents two novel low-computational pattern recognition approaches for identifying abnormal conveyor belt usage in real-time on resource-constrained devices. The performance is evaluated and compared with a former method. (ii) The approaches are tested for machine drift over time on an additional dataset acquired more than a year after training. (iii) The approaches are optimized using quantization and implemented on two low-power MCUs, measuring energy consumption and memory usage.
@@ II Conveyor Belt Operation
A hydraulic conveyor belt system operates by harnessing fluid power to drive the movement of the belt through hydraulic components. It is a highly efficient continuous transport mechanism, capable of moving heavy loads over long distances. In this system, hydraulic charge pumps pressurize fluid, which is directed through control valves. These valves regulate the flow and direction of oil to drive the hydraulic motors attached to the drive pulleys or rollers of the conveyor belt. The resulting rotational motion propels the belt, facilitating the seamless transportation of materials along the designated path. The control system provides flexibility in adjusting the speed and direction of the belt. It also offers precise control over its operation. Additionally, sensors are connected to the control system. These sensors continuously track and log key hydraulic parameters.
@@ III-A Datasets
This study utilized two datasets, each derived from parameters monitored by the control system of a conveyor belt. The conveyor control system maintains a fixed belt speed regardless of material load. Variations in material load alter the required motor torque, causing proportional changes in hydraulic pressure. Monitoring these pressure fluctuations enables real-time assessment of conveyor workload and system stress. Two pressure sensors recorded the high- and low-pressure of the hydraulic motor in bar. Additionally, an incremental encoder with a pulsed output measured the number of motor revolutions per minute (rpm). These sensors were integrated into the control system through a 4-20 mA interface. The collected sensor data underwent preprocessing to achieve a sampling rate of one measurement per minute. These values were retrieved from the control system’s user interface. For each sensor, the one-minute interval values represent the average of all readings collected at the sensor’s nominal sampling frequency.
The first dataset (DS1) comprises data from four months representing all four seasons: June 2021, October 2021, January 2022, and April 2022. This temporal coverage captured a broad temperature range (-26.1 to 26.0°C), significantly affecting oil viscosity and consequently the pressure levels. The second dataset (DS2) was collected from the same conveyor belt during June, August, October, and December 2023, with temperatures ranging from -29.0 to 26.0°C. Figure 1 shows the distribution of the three sensed variables in DS1 and DS2: high-pressure, low-pressure, and speed. Whisker values in the box plots represent the minimum and maximum values. Both datasets show similar spreads for most high-pressure values, though DS2 covers a wider range. DS1’s low-pressure values cover a greater range than DS2’s, with notable quartile differences. This difference could be caused by machine wear and maintenance. Speed values cluster around 0 and 45 rpm in both datasets when the conveyor belt is stopped or moving, respectively. Variations between these values occur during belt acceleration or deceleration.
TABLE I: Operation modes samples in DS1.
Samples (percentage) Class
10,566 (6.2%) Off
123,225 (72.7%) Idle
11,454 (6.8%) Operational
24,299 (14.3%) Active
TABLE II: Duty cycle samples distribution in DS1 and DS2.
Dataset Samples (percentage) Class
DS1 496 (82.3%) Normal
107 (17.7%) Abnormal
DS2 751 (79.5%) Normal
194 (20.5%) Abnormal
DS1 contains approximately 170,000 operation mode samples in 1-minute intervals, labeled shown in Table I. Table II presents the number of normal and abnormal duty cycles in each dataset. DS1 was used to develop and evaluate the proposed approaches, while DS2 was used to assess the generalization and robustness of the approaches in a different dataset.
Hydraulic system experts analyzed the waveform patterns using the Imagimob Studio software (Imagimob AB, Stockholm, Sweden) to determine the start and end timestamps, and the ground truth labels for the operation modes and duty cycles.
@@ III-B Algorithms
This study compares a state-of-the-art and two novel approaches to identify anomalies in conveyor belt operations. The approaches classify the duty cycle usages of a conveyor belt applying pattern recognition techniques.
@@
    Approach-1 proposed by Martinez-Rau et al. [17], consists of four consecutive stages (Fig. 2(a)) and serves as a baseline. The algorithm takes as input the three 1-min sensed variables (motor shaft rotational speed, and the high- and low-pressure of the hydraulic motor), which are used to extract a set of 12 features. Four features consist of the three input values plus the differential pressure between the high- and low-pressure. The remaining eight features are derived from the former four features. These are third- and fifth-order moving averages, which effectively capture fast and slow variations in the input variables. The complete feature set feeds an ML model to classify the operation modes labeled every 1 min. Since variations in the operation modes typically occur at lower frequencies, a third-order median filter smooths the predicted labels to reduce label fragmentation. Finally, transitions in the filtered operation mode labels are analyzed through pattern matching to detect and classify normal and abnormal duty cycles. Therefore, the correct detection and classification of duty cycles depends on the correct prediction of the operation modes.
@@
    Approach-2 offers an alternative method for detecting duty cycles (Fig. 2(b)). Approach-1 relies on accurately classifying operation mode labels to detect and classify duty cycles. In contrast, Approach-2 bypasses operation mode classification and focuses directly on input speed values. Based on the speed values distribution (Fig. 1), an empirical threshold of 5 rpm is chosen to determine the boundaries of the duty cycles. Once these boundaries are identified, the remaining steps in Approach-2 follow a similar process to those in Approach-1.
@@
    Approach-3 builds on the concept of Approach-2 but introduces an ML model to classify duty cycles instead of relying on the pattern-matching rules used in the previous approaches (Fig. 2(c)). By employing an ML-based classifier, Approach-3 offers the advantage of generating smoother decision boundaries, which can help mitigate the impact of potential misclassifications of the operation modes. In this approach, each transition mode label links to an input of the duty-cycle classifier. However, since the number of transitions within a duty cycle can vary, a value of 20 inputs was defined arbitrarily. This value exceeds the maximum observed transitions (15) in both datasets. Entries not associated with a transition mode label are assigned a different label.
The operation mode and the duty-cycle classifiers are evaluated with different combination of supervised ML algorithms: DT, RF, ET, XGB, NB, and MLP.
III-C Experimental Setup
The experiments were implemented in Python 3.8.10, utilizing the sklearn 1.2.2 package for tree-based and NB algorithms, and TensorFlow 2.13.0 for the MLP.
TABLE III: Main properties of the MCU platforms
MCU Processor Test board CPU frequency SRAM Flash FPU Current consumption Voltage
@@ III-C Experimental Setup
The three approaches were trained and evaluated in the first experiment using DS1. The operation mode classifiers were trained and tested using a leave-one-month-out cross-validation strategy. In each iteration, three months of data were used for training, while the remaining month served for testing. Synthetic oversampling and random undersampling techniques were applied using the SMOTE algorithm to balance the training set [24]. In each leave-one-month-out iteration, 100% of the minority Off and Operational classes were added in the training set, while 80% of the majority Idle classes were removed. The gridsearchCV method of sklearn was used for hyperparameter tuning to optimize model performance. This method applied 5-fold cross-validation on the balanced training set to train and evaluate the model with fixed hyperparameters. The macro-averaged F1-score per class was computed in each fold. Then, a final fold-averaged F1-score was computed to select the best hyperparameter values. The following hyperparameters were evaluated during the optimization. The maximum depth of the DT was set with a restriction of 10, 20, 30, 40, and 50 splits, as well as without restriction. In all cases, the Gini impurity and the Shannon entropy were considered as criteria for the quality split [21]. For the tree-based ensemble algorithms, the number of trees was set to 10, 25, and 50, with maximum tree depths of 4, 6, 8, and 10. The MLP architecture employed a single hidden layer, with the number of neurons varying between 4 and 15. The number of neurons in the input layer was fixed at 12 (number of features) and the number of neurons in the output layer was 4 (number of operation mode labels). Training was conducted using stochastic gradient descent, with learning rates of 0.1, 0.01, and 0.001. The other hyperparameters kept the default values defined in the respective Python packages.
In Approach-1 and Approach-2, the predicted operation mode labels were used to detect and classify duty cycles based on heuristic rules (Fig. 2). Similar to the operation mode classifier, the duty-cycle classifier in Approach-3 was trained and tested using leave-one-month-out cross-validation. In each iteration, the number of abnormal duty cycles was augmented using the SMOTEN algorithm to have the same number of normal and abnormal cycle samples in the training set [24]. Hyperparameter optimization was performed in the same way as when training the operation mode classifier. The operation mode classifier was trained first, then the duty-cycle classifier was trained and evaluated using the predicted operation mode labels across the four months. This approach allows the duty-cycle classifier to learn and adapt to potential misclassifications made by the operation mode classifier.
The generalization capabilities of the proposed approaches were evaluated in the second experiment, where the operation mode and duty-cycle classifiers were trained on the entire DS1 dataset and tested on DS2. During training, the same hyperparameter optimization and class balancing techniques were applied as in the first experiment.
A key factor to consider was the real-time operational requirements for anomaly monitoring in conveyor belts. To assess resource-constrained hardware requirements, the best-performing models in each dataset were implemented in C code. The best-performing ML models from the previous two experiments were converted to C code using the emlearn 0.19.3 package [^1] https://github.com/emlearn/emlearn/tree/0.19.3 for tree-based and NB models, and LiteRT [^2] https://ai.google.dev/edge/litert for the MLP. Similar to the second experiment, the models were trained on DS1 and tested on DS2. The energy consumption, memory usage, and algorithm performance were analyzed using ML models with 32-bit float resolutions and post-training 8-bit integer full model quantization. The compiled C code, using arm-none-eabi-g++, was deployed on two MCU boards: the Raspberry Pi Pico and the Arduino Nano Sense 33 BLE Lite. Table III provides the specifications of these boards.
@@
where $TP_{i}$ counts the number of correctly classified instances of class $i$, $FP_{i}$ counts the number of recognized instances of class $i$ that do not exist in the ground truth (insertions), and $FN_{i}$ counts the number of ground truth instances of class $i$ that are not recognized (deletions). An overall micro-average F1-score was also computed to aggregate the total false positives, false negatives, and true positives across both classes [26]. Except for the deployment on MCUs, each train/test experiment was repeated with different random seeds ten times to ensure robustness. For Approach-3, 100 iterations were performed by combining the ten repetitions of the duty-cycle classifiers with the ten repetitions of the operation mode classifiers. The final metrics reported are the averages over all repetitions.